4. Assume a wheat producing farmer engaging in selling its product under perfect competition market faces cost functions as TC= Q3 -2Q2 +8Q and Average revenue of the farmer is given as Birr 8 . Having this information, A) Determine the optimal level of output and price in the short run. B) Calculate the economic profit (loss) the farmer will obtain (incur) C) What will be the minimum price level the farmer gets to continue in wheat production


1
Expert's answer
2022-01-04T10:20:42-0500

4.

A) The optimal level of output and price in the short run is at P = MC = MR = AR.

"MC = TC'(Q) = 3Q^2 - 4Q + 8 = 8,"

"3Q^2 - 4Q = 0,"

Q = 0 (not acceptable) or Q = 4/3 units.

P = 8.


B) The economic profit the farmer will obtain is:

"TP = 8\u00d74\/3 - ((4\/3)^3 - 2\u00d7(4\/3)^2 + 8\u00d74\/3) = 1.19."


C) The minimum price level the farmer gets to continue in wheat production is at P = AVC.